OpenStudy (anonymous):

-9sqrt32 + 9sqrt18 +10sqrt8

OpenStudy (lgbasallote):

\[-9\sqrt {16 \times 2} + 9 \sqrt{9 \times 2} + 10 \sqrt{4\times 2}\] \[-9(4\sqrt 2) + 9(3\sqrt 2) + 10(2\sqrt 2)\]

OpenStudy (lgbasallote):

do you get it?

OpenStudy (anonymous):

-36sqrt2 27sqrt2 20sqrt2

OpenStudy (lgbasallote):

\[-36\sqrt 2 + 27\sqrt 2 + 20\sqrt 2\] right

OpenStudy (lgbasallote):

now combine

OpenStudy (anonymous):

11sqrt2

OpenStudy (lgbasallote):

correct

OpenStudy (anonymous):

Thanks!

OpenStudy (lgbasallote):

<tips hat>